# The separating Noether number of small groups

### v2

@@ 6.1. The group $\mathrm{Dic}_{8}\times\mathrm{C}_{2}$
By (2) we have $\beta_{\mathrm{sep}}^{K}(G)\geq\beta_{\mathrm{sep}}^{K}(\mathrm{A}_{4})$, and by Proposition 5.1 we have $\beta_{\mathrm{sep}}^{K}(\mathrm{A}_{4})\geq 6$. By Lemma 3.1 (iii), to prove the reverse inequality $\beta_{\mathrm{sep}}^{K}(G)\leq 6$ it is sufficient to show that $\beta_{\mathrm{sep}}^{L}(G)\leq 6$ for some extension $L$ of $K$. Therefore we may assume that $K$ is large enough to apply Lemma 3.6, and so it is sufficient to prove $\beta_{\mathrm{sep}}(G,V)\leq 6$ for $V:=W_{1}\oplus W_{2}\oplus U$. Take $v=(w_{1},w_{2},u)\in V$, $v^{\prime}=(w^{\prime}_{1},w^{\prime}_{2},u^{\prime})\in V$, and assume that $f(v)=f(v^{\prime})$ for all $f\in K[V]^{G}_{\leq 6}$. We need to show that then $G\cdot v=G\cdot v^{\prime}$. By Proposition 6.12 we know that $G\cdot(w_{1},w_{2})=G\cdot(w^{\prime}_{1},w^{\prime}_{2})$, so replacing $v^{\prime}$ by an appropriate element in its $G$-orbit we may assume that $w^{\prime}_{1}=w_{1}$, $w^{\prime}_{2}=w_{2}$. By Proposition 6.10 (ii) it is sufficient to deal with the case $w_{1}\neq 0$. The Davenport constant of $G/G^{\prime}$ is $6$, therefore $G\cdot u=G\cdot u^{\prime}$. By Proposition 6.12 it is sufficient to deal with the case when $u\neq 0$, moreover, when $u$ has a non-zero component $u_{\chi}$ for some non-identity character $\chi\in\widehat{G}$.
@@
Case II.b: $|\mathrm{Stab}_{G}(w_{1})|=3$ (consequently, by Lemma 6.11 we have $x_{1}^{2}(v)=x_{2}^{2}(v)=x_{3}^{2}(v)\neq 0$, hence in particular, $(x_{1}x_{2}x_{3})(v)\neq 0$) and $\mathrm{Stab}_{G}(w_{1})\subseteq\mathrm{Stab}_{G}(w_{2})$. Write $H:=\mathrm{Stab}_{G}(w_{1})$. Then $H\cong HG^{\prime}/G^{\prime}$ (since $|H|=3$ is coprime to $|G^{\prime}|=4$). It follows that an $H$-invariant monomial $h$ in $K[U]$ is either $G$-invariant or is a relative $G$-invariant with weight $(1,-1)$. Hence $h$ or $x_{1}x_{2}x_{3}h$ is a $G$-invariant, implying by $(x_{2}x_{2}x_{3})(v)\neq 0$ that $h(u)=h(u^{\prime})$ holds for all $H$-invariant monomials in $K[U]$ of degree at most $3=\mathsf{D}(\mathrm{C}_{3})=\mathsf{D}(H)=\beta_{\mathrm{sep}}^{K}(H)$. Consequently, there exists an element $g\in H$ with $g\cdot u=u^{\prime}$. Now $g\cdot(w_{1},w_{2},u)=(w_{1},w_{2},u^{\prime})$, so $G\cdot v=G\cdot v^{\prime}$.
Case II.c: $|\mathrm{Stab}_{G}(w_{1})|=3$ (consequently, by Lemma 6.11 we have $x_{1}^{2}(v)=x_{2}^{2}(v)=x_{3}^{2}(v)\neq 0$) and $y_{1}^{2}(v)=y_{2}^{2}(v)=y_{2}^{2}(v)\neq 0$. Replacing the pair $(v,v^{\prime})$ by an appropriate element in its $G$-orbit we may assume that $w_{1}$ is a non-zero scalar multiple of $[1,1,1]^{T}$. If $w_{2}$ is also a scalar multiple of $[1,1,1]^{T}$, then we are in Case II.b. So we may assume that $w_{2}$ is not a scalar multiple of $[1,1,1]^{T}$. Then necessarily $w_{2}$ is a non-zero scalar multiple of one of $[-1,1,1]^{T}$, $[1,-1,1]^{T}$, $[1,1,-1]^{T}$. Then the relative invariants
@@
Case III.b: $y_{1}(v)=0$ and $(y_{2}y_{3})(v)=0$. If $y_{3}(v)=0$, then $H:=\langle ad\rangle\subseteq\mathrm{Stab}_{G}(w_{1},w_{2})$. Note that $H\cong HG^{\prime}/G^{\prime}$. A monomial $m\in K[U]^{H}$ is a relative $G$-invariant with weight in $\langle(\omega,1)\rangle$. If $\deg(m)\leq 2$, then by Lemma 6.13 (iii) we have a $G$-invariant $mf$ with $\deg(mf)\leq 6$ and $f(w_{1},w_{2})\neq 0$. It follows that $m(u)=m(u^{\prime})$ for all $H$-invariant monomials $m\in K[U]$ with $\deg(m)\leq 2=\mathsf{D}(H)=\beta_{\mathrm{sep}}(H)$. Consequently, $H\cdot u=H\cdot u^{\prime}$, and as $H$ stabilizes $(w_{1},w_{2})$, we conclude $G\cdot v=G\cdot v^{\prime}$. The case $y_{2}(v)=0$ is similar, we just need to take $H:=\langle abd\rangle\subseteq\mathrm{Stab}_{G}(w_{1},w_{2})$ in the above argument.
@@
Case IV.b: exactly one of $y_{1}(v)$, $y_{2}(v)$ is zero. Then $h:=x_{1}y_{1}+x_{2}y_{2}+x_{3}y_{3}$ is a degree $\leq 3$ relative $G$-invariant with weight $(1,-1)$ and not vanishing at $(w_{1},w_{2})$, so we are done, as we explained in the beginning of Case IV.
@@
in $K[W_{1}\oplus W_{2}]^{G,\chi}$. One of them does not vanish at $(w_{1},w_{2})$. Indeed, suppose for contradiction that all of them vanish at $(w_{1},w_{2})$. From $r_{(\omega,-1)}^{(1)}(w_{1},w_{2})=0=r_{(\omega,-1)}^{(2)}(w_{1},w_{2})$ we deduce $y_{1}^{2}(w_{2})=y_{2}^{2}(w_{2})$. Thus $[y_{1}(v),y_{2}(v)]$ is a non-zero scalar multiple of $[1,1]$ or $[1,-1]$. After that, from $r_{(\omega,-1)}^{(1)}(w_{1},w_{2})=0=r_{(\omega,-1)}^{(3)}(w_{1},w_{2})$ we conclude $x_{1}^{2}(v)=x_{2}^{2}(v)$ and hence $x_{1}(v)=\pm x_{2}(v)$, which clearly contradicts to the assumption that $r_{(\omega,-1)}^{(1)}(w_{1},w_{2})=0$. The case of the weight $\chi=(\omega^{2},-1)$ can be settled similarly, using the relative invariants $x_{1}y_{1}+\omega^{2}x_{2}y_{2}+\omega x_{3}y_{3}$, $x_{1}y_{1}^{3}+\omega^{2}x_{2}y_{2}^{3}+\omega x_{3}y_{3}^{3}$, $x_{1}^{3}y_{1}+\omega^{2}x_{2}^{3}y_{2}+\omega x_{3}^{3}y_{3}$. ∎
@@ 7. The group $\mathrm{S}_{3}\times\mathrm{C}_{3}$
Assume in addition that $K$ contains an element $\omega$ of multiplicative order $3$, and consider the following irreducible $2$-dimensional representations of $G$:
@@
For example, $q_{xx}=x_{1}x_{2}$, $q_{yz}=\frac{1}{2}(y_{1}z_{2}+y_{2}z_{1})$, $p_{xxx}=x_{1}^{3}+x_{2}^{3}$, $p_{xxy}=x_{1}^{2}y_{1}+x_{2}^{2}y_{2}$. It is well known (see for example [33, Theorem 4.1]) that the elements $q_{\xi\eta}$, $p_{\xi\eta\zeta}$ minimally generate $K[W_{1}\oplus W_{2}\oplus W_{3}]^{\mathrm{S}_{3}}$, where $\mathrm{S}_{3}$ is identified with the subgroup $\langle a,b\rangle$ of $G$. Set
$A_{(1,1)}$ :=$:=\{q_{xx},q_{yz},p_{xxx},p_{yyy},p_{zzz},p_{xyz}\}$
$A_{(1,\omega)}$ :=$:=\{q_{xy},q_{zz},p_{xxy},p_{xzz},p_{yyz}\}$
$A_{(1,\omega^{2})}$ :=$:=\{q_{xz},q_{yy},p_{xxz},p_{xyy},p_{yzz}\}.$
@@
By [15, Proposition 3.3] we have the inequality $\beta^{K}(G)=8$, so we can omit the degree $9$ elements from the generating system $A$. Therefore in order to prove that $B$ generates $K[W_{1}\oplus W_{2}\oplus W_{3}]^{G}$, it is sufficient to show that all the elements of $A\setminus B$ of degree $7$ or $8$ are contained in the ideal of $K[W_{1}\oplus W_{2}\oplus W_{3}]$ generated by the invariants of degree at most $6$. All elements of degree $7$ or $8$ in $A\setminus B$ are products of the form $f_{1}f_{2}f_{3}$ where $\deg(f_{1})=2$, $\deg(f_{2})=3$, and both of $f_{1}$ and $f_{2}$ belong either to $A_{(1,\omega)}$ or to $A_{(1,\omega^{2})}$. The equalities
@@
By assumption we have
$(q_{zz}q_{xz})(v)=(q_{zz}q_{xz})(v^{\prime}).$
hence
@@
Again by assumption we have
$(p_{xzz}q_{xz})(v)=(p_{xzz}q_{xz})(v^{\prime}).$
Therefore if $q_{xz}(v)\neq 0$, then we have
@@
Consequently, $w^{\prime}_{1}$ is a non-zero scalar multiple of $w_{1}$. On the other hand, $w^{\prime}_{1}\in G\cdot w_{1}$, thus $w_{1}$ is an eigenvector of some matrix in $\{\psi_{1}(g)\mid g\in G\}=\psi_{1}(G)$. If the corresponding eigenvalue is $1$, then $(w_{1},w_{3})=(w^{\prime}_{1},w^{\prime}_{3})$ and (21) holds. If the eigenvalue is not $1$, then $w_{1}$ is the $-1$-eigenvector of some non-diagonal element of $\psi_{1}(G)$ (since the eigenvectors with eigenvalue different from $1$ of the diagonal elements in $\psi_{1}(G)$ are scalar multiples of $[1,0]^{T}$ or $[0,1]^{T}$, and $w_{1}$ is not of this form by (24) and (23)). After rescaling $W_{1}$ and $W_{3}$ (see Section 3.3), we have that
@@
$(w_{1},w_{3})=([\omega^{2},-1]^{T},[\omega^{2},1]^{T})\text{ and }(w^{\prime}_{1},w^{\prime}_{3})=([-\omega^{2},1]^{T},[\omega^{2},1]^{T})$
@@
We have $K[W_{1}]^{G,(-1,1)}=Kp^{-}_{xxx}\oplus(\mathcal{H}(G,W_{1})\cap K[W_{1}]^{G,(-1,1)})$, and $K[W_{1}]^{G,\chi}=\{0\}$ for all $\chi\in\widehat{G}\setminus\{(\pm 1,1)\}$. Therefore by Lemma 3.9, $K[W_{1}\oplus U]^{G}$ is generated by $K[W_{1}]^{G}$, $K[U]^{G}$, and
@@
Since $\mathsf{D}(\mathrm{C}_{6}/\mathrm{C}_{2})=\mathsf{D}(\mathrm{C}_{3})=3$, for $k\geq 3$ the sequence $\chi^{(1)},\dots\chi^{(k)}$ necessarily has a subsequence with product (±$(\pm 1,1)$, so the above generators of the form $p^{-}_{xxx}t_{\chi^{(1)}}\cdots t_{\chi^{(k)}}$ have degree at most $3+3$. We have $K[W_{1}]^{G}=K[q_{xx},p_{xxx}]$, and $\beta(G,U)=\mathsf{D}(\mathrm{C}_{6})=6$. Thus each generator of $K[W_{1}\oplus U]^{G}$ has degree at most $6$. ∎
@@
The case $W=W_{3}$ follows by Lemma 3.11. Indeed, the group $G$ has the automorphism $\alpha:a\mapsto a$, $b\mapsto b$, $c\mapsto c^{-1}$. We have $\psi_{2}=\psi_{3}\circ\alpha$, so $\alpha$ interchanges $W_{2}$ and $W_{3}$, and permutes the $1$-dimensional representations. ∎
@@
Case 1.a: $q_{yy}(w_{2})=0=q_{zz}(w_{3})$. Then both $w_{2}$ and $w_{3}$ are non-zero scalar multiples of $[1,0]^{T}$ or $[0,1]^{T}$. There is no harm to assume that $w_{2}=[1,0]^{T}$ (the case $w_{2}\in K^{\times}[0,1]^{T}$ is similar). Then $\mathrm{Stab}_{G}(w_{2})=\langle ac^{2}\rangle$, and $w_{3}$ can not be a scalar multiple of $[0,1]^{T}$ (since the stabilizer of $[0,1]^{T}\in W_{3}$ is also $\langle ac^{2}\rangle$). This means that $w_{3}$ is a non-zero scalar multiple of $[1,0]^{T}$ as well, and so none of $\{p_{yyz},p_{yzz},p^{-}_{yyz},p^{-}_{yzz},p^{-}_{zzz}\}$ vanishes on $v=(w_{2},w_{3})$. These are relative invariants of degree $3$, and their weights represent all the five non-trivial characters.
Case 1.b: $q_{yy}(w_{2})=0=p^{-}_{zzz}(w_{3})$. Then $w_{2}$ is a non-zero scalar multiple of $[1,0]^{T}$ or $[0,1]^{T}$, and $w_{3}$ is a non-zero scalar multiple of $[1,1]^{T}$, $[1,\omega]^{T}$, $[1,\omega^{2}]^{T}$. It follows that none of $\{p_{yyz},p_{yzz},p^{-}_{yyz},p^{-}_{yzz},p^{-}_{yyy}\}$ vanishes on $v$. These are relative invariants of degree $3$, and their weights represent all the five non-trivial characters.
@@
Case 1.d: $p^{-}_{yyy}(w_{2})=0=p^{-}_{zzz}(w_{3})$. Then both of $w_{2},w_{3}$ are non-zero scalar multiples of $[1,1]^{T}$, $[1,\omega]^{T}$, or $[1,\omega^{2}]^{T}$. Moreover, $w_{2}$ and $w_{3}$ are linearly independent (otherwise they would have the same stabilizer). Thus none of $\{q_{yy},q_{zz},q^{-}_{yz},q_{yy}q^{-}_{yz},q_{zz}q^{-}_{yz}\}$ vanishes on $v$. These are relative invariants of degree at most $4$, and their weights represent all the five non-trivial characters.

### v3

@@ 4. Some groups $G=H\times\mathrm{C}_{2}$ with $\beta_{\mathrm{sep}}^{K}(G)=\beta_{\mathrm{sep}}^{K}(H)$
Since $\mathrm{A}_{4}$ is a homomorphic image of $G$, we have $\beta_{\mathrm{sep}}^{K}(G)\geq\beta_{\mathrm{sep}}^{K}(\mathrm{A}_{4})$, and by [15, Proposition 5.1] we have $\beta_{\mathrm{sep}}^{K}(\mathrm{A}_{4})\geq 6$. By Lemma 2.3 (iii), to prove the reverse inequality $\beta_{\mathrm{sep}}^{K}(G)\leq 6$ it is sufficient to show that $\beta_{\mathrm{sep}}^{L}(G)\leq 6$ for some extension $L$ of $K$. Therefore we may assume that $K$ is large enough to apply Lemma 2.4, and so it is sufficient to prove $\beta_{\mathrm{sep}}(G,V)\leq 6$ for $V:=W_{1}\oplus W_{2}\oplus U$. Take $v=(w_{1},w_{2},u)\in V$, $v^{\prime}=(w^{\prime}_{1},w^{\prime}_{2},u^{\prime})\in V$, and assume that $f(v)=f(v^{\prime})$ for all $f\in K[V]^{G}_{\leq 6}$. We need to show that then $G\cdot v=G\cdot v^{\prime}$. By Proposition 4.8 we know that $G\cdot(w_{1},w_{2})=G\cdot(w^{\prime}_{1},w^{\prime}_{2})$, so replacing $v^{\prime}$ by an appropriate element in its $G$-orbit we may assume that $w^{\prime}_{1}=w_{1}$, $w^{\prime}_{2}=w_{2}$. By Proposition 4.6 (ii) it is sufficient to deal with the case $w_{1}\neq 0$. The Davenport constant of $G/G^{\prime}$ is $6$, therefore $G\cdot u=G\cdot u^{\prime}$. By Proposition 4.8 it is sufficient to deal with the case when $u\neq 0$, moreover, when $u$ has a non-zero component $u_{\chi}$ for some non-identity character $\chi\in\widehat{G}$.
@@
Case II.b: $|\mathrm{Stab}_{G}(w_{1})|=3$ (consequently, by Lemma 4.7 we have $x_{1}^{2}(v)=x_{2}^{2}(v)=x_{3}^{2}(v)\neq 0$, hence in particular, $(x_{1}x_{2}x_{3})(v)\neq 0$) and $\mathrm{Stab}_{G}(w_{1})\subseteq\mathrm{Stab}_{G}(w_{2})$. Write $H:=\mathrm{Stab}_{G}(w_{1})$. Then $H\cong HG^{\prime}/G^{\prime}$ (since $|H|=3$ is coprime to $|G^{\prime}|=4$). It follows that an $H$-invariant monomial $h$ in $K[U]$ is either $G$-invariant or is a relative $G$-invariant with weight (1,1)(1,-1). Hence $h$ or $x_{1}x_{2}x_{3}h$ is a $G$-invariant, implying by $(x_{1}x_{2}x_{3})(v)\neq 0$ that $h(u)=h(u^{\prime})$ holds for all $H$-invariant monomials in $K[U]$ of degree at most $3=\mathsf{D}(\mathrm{C}_{3})=\mathsf{D}(H)=\beta_{\mathrm{sep}}^{K}(H)$. Consequently, there exists an element $g\in H$ with $g\cdot u=u^{\prime}$. Now $g\cdot(w_{1},w_{2},u)=(w_{1},w_{2},u^{\prime})$, so $G\cdot v=G\cdot v^{\prime}$.
Case II.c: $|\mathrm{Stab}_{G}(w_{1})|=3$ (consequently, by Lemma 4.7 we have $x_{1}^{2}(v)=x_{2}^{2}(v)=x_{3}^{2}(v)\neq 0$) and $y_{1}^{2}(v)=y_{2}^{2}(v)=y_{3}^{2}(v)\neq 0$. Replacing the pair $(v,v^{\prime})$ by an appropriate element in its $G$-orbit we may assume that $w_{1}$ is a non-zero scalar multiple of $1,1,1]^{T}$. If $w_{2}$ is also a scalar multiple of $1,1,1]^{T}$, then we are in Case II.b. So we may assume that $w_{2}$ is not a scalar multiple of $1,1,1]^{T}$. Then necessarily $w_{2}$ is a non-zero scalar multiple of one of $1,1]^{T}, [1,1,$$1]^{T}, [1,1,$$1,-1]^{T}$. Then the relative invariants
@@
Case III.b: $y_{1}(v)=0$ and $(y_{2}y_{3})(v)=0$. If $y_{3}(v)=0$, then $H:=\langle ad\rangle\subseteq\mathrm{Stab}_{G}(w_{1},w_{2})$. Note that $H\cong HG^{\prime}/G^{\prime}$. A monomial $m\in K[U]^{H}$ is a relative $G$-invariant with weight in $\langle(\omega,1)\rangle$. If $\deg(m)\leq 2$, then by Lemma 4.9 (iii) we have a $G$-invariant mfmf with $\deg(mf)\leq 6$ and $f(w_{1},w_{2})\neq 0$. It follows that $m(u)=m(u^{\prime})$ for all $H$-invariant monomials $m\in K[U]$ with $\deg(m)\leq 2=\mathsf{D}(H)=\beta_{\mathrm{sep}}(H)$. Consequently, $H\cdot u=H\cdot u^{\prime}$, and as $H$ stabilizes $(w_{1},w_{2})$, we conclude $G\cdot v=G\cdot v^{\prime}$. The case $y_{2}(v)=0$ is similar, we just need to take $H:=\langle abd\rangle\subseteq\mathrm{Stab}_{G}(w_{1},w_{2})$ in the above argument.
@@
Case IV.b: exactly one of $y_{1}(v)$, $y_{2}(v)$ is zero. Then $h:=x_{1}y_{1}+x_{2}y_{2}+x_{3}y_{3}$ is a degree $\leq 3$ relative $G$-invariant with weight (1,1)(1,-1) and not vanishing at $(w_{1},w_{2})$, so we are done, as we explained in the beginning of Case IV.
@@
in $K[W_{1}\oplus W_{2}]^{G,\chi}$. One of them does not vanish at $(w_{1},w_{2})$. Indeed, suppose for contradiction that all of them vanish at $(w_{1},w_{2})$. From $r_{(\omega,-1)}^{(1)}(w_{1},w_{2})=0=r_{(\omega,-1)}^{(2)}(w_{1},w_{2})$ we deduce $y_{1}^{2}(w_{2})=y_{2}^{2}(w_{2})$. Thus $y_{1}(v),y_{2}(v)]$ is a non-zero scalar multiple of $[1,1]$ or [1,1][1,-1]. After that, from $r_{(\omega,-1)}^{(1)}(w_{1},w_{2})=0=r_{(\omega,-1)}^{(3)}(w_{1},w_{2})$ we conclude $x_{1}^{2}(v)=x_{2}^{2}(v)$ and hence $x_{1}(v)=\pm x_{2}(v)$, which clearly contradicts to the assumption that $r_{(\omega,-1)}^{(1)}(w_{1},w_{2})=0$. The case of the weight $\chi=(\omega^{2},-1)$ can be settled similarly, using the relative invariants $x_{1}y_{1}+\omega^{2}x_{2}y_{2}+\omega x_{3}y_{3}$, $x_{1}y_{1}^{3}+\omega^{2}x_{2}y_{2}^{3}+\omega x_{3}y_{3}^{3}$, $x_{1}^{3}y_{1}+\omega^{2}x_{2}^{3}y_{2}+\omega x_{3}^{3}y_{3}$. ∎
@@ 5. The group $\mathrm{S}_{3}\times\mathrm{C}_{3}$
Assume in addition that $K$ contains an element $\omega$ of multiplicative order $3$, and consider the following pairwise non-isomorphic irreducible $2$-dimensional representations of $G$:
@@
For example, $q_{xx}=x_{1}x_{2}$, $q_{yz}=\frac{1}{2}(y_{1}z_{2}+y_{2}z_{1})$, $p_{xxx}=x_{1}^{3}+x_{2}^{3}$, $p_{xxy}=x_{1}^{2}y_{1}+x_{2}^{2}y_{2}$. It is well known (see for example [22, Theorem 4.1]) that the elements $q_{\xi\eta}$, $p_{\xi\eta\zeta}$ minimally generate $K[W_{1}\oplus W_{2}\oplus W_{3}]^{\mathrm{S}_{3}}$, where $\mathrm{S}_{3}$ is identified with the subgroup $\langle a,b\rangle$ of $G$. Set
$A_{(1,1)}$ $:=\{q_{xx},q_{yz},p_{xxx},p_{yyy},p_{zzz},p_{xyz}\}$
$A_{(1,\omega)}$ $:=\{q_{xy},q_{zz},p_{xxy},p_{xzz},p_{yyz}\}$
$A_{(1,\omega^{2})}$ $:=\{q_{xz},q_{yy},p_{xxz},p_{xyy},p_{yzz}\}.$
@@
By [11, Proposition 3.3] we have the inequality $\beta^{K}(G)=8$, so we can omit the degree $9$ elements from the generating system $A$. Therefore in order to prove that $B$ generates $K[W_{1}\oplus W_{2}\oplus W_{3}]^{G}$, it is sufficient to show that all the elements of $A\setminus B$ of degree $7$ or $8$ are contained in the ideal of $K[W_{1}\oplus W_{2}\oplus W_{3}]$ generated by the invariants of degree at most $6$. All elements of degree $7$ or $8$ in $A\setminus B$ are products of the form $f_{1}f_{2}f_{3}$ where $\deg(f_{1})=2$, $\deg(f_{2})=3$, and both of $f_{1}$ and $f_{2}$ belong either to $A_{(1,\omega)}$ or to $A_{(1,\omega^{2})}$. The equalities
@@
By assumption we have $(q_{zz}q_{xz})(v)=(q_{zz}q_{xz})(v^{\prime})$, hence
@@
Again by assumption we have $(p_{xzz}q_{xz})(v)=(p_{xzz}q_{xz})(v^{\prime})$. Therefore if $q_{xz}(v)\neq 0$, then we have
@@
Consequently, $w^{\prime}_{1}$ is a non-zero scalar multiple of $w_{1}$. On the other hand, $w^{\prime}_{1}\in G\cdot w_{1}$, thus $w_{1}$ is an eigenvector of some matrix in $\{\psi_{1}(g)\mid g\in G\}=\psi_{1}(G)$. If the corresponding eigenvalue is $1$, then $(w_{1},w_{3})=(w^{\prime}_{1},w^{\prime}_{3})$ and (11) holds. If the eigenvalue is not $1$, then $w_{1}$ is the 1-1-eigenvector of some non-diagonal element of $\psi_{1}(G)$ (since the eigenvectors with eigenvalue different from $1$ of the diagonal elements in $\psi_{1}(G)$ are scalar multiples of $1,0]^{T}$ or $0,1]^{T}$, and $w_{1}$ is not of this form by (14) and (13)). After rescaling $W_{1}$ and $W_{3}$ (see Section 2.1, (3)), we have that
@@
[…]
@@
[…]
@@
Since $\mathsf{D}(\mathrm{C}_{6}/\mathrm{C}_{2})=\mathsf{D}(\mathrm{C}_{3})=3$, for $k\geq 3$ the sequence $\chi^{(1)},\dots\chi^{(k)}$ necessarily has a subsequence with product $(\pm 1,1)$, so the above generators of the form $p^{-}_{xxx}t_{\chi^{(1)}}\cdots t_{\chi^{(k)}}$ have degree at most $3+3$. We have $K[W_{1}]^{G}=K[q_{xx},p_{xxx}]$, and $\beta(G,U)=\mathsf{D}(\mathrm{C}_{6})=6$. Thus each generator of $K[W_{1}\oplus U]^{G}$ has degree at most $6$. ∎
@@
The case $W=W_{3}$ follows by Lemma 2.5 (ii). Indeed, the group $G$ has the automorphism $\alpha:a\mapsto a$, $b\mapsto b$, $c\mapsto c^{-1}$. We have $\psi_{2}=\psi_{3}\circ\alpha$, so $\alpha$ interchanges $W_{2}$ and $W_{3}$, and permutes the $1$-dimensional representations. ∎
@@
Case 1.a: $q_{yy}(w_{2})=0=q_{zz}(w_{3})$. Then both $w_{2}$ and $w_{3}$ are non-zero scalar multiples of $1,0]^{T}$ or $0,1]^{T}$. There is no harm to assume that $w_{2}=[1,0]^{T}$ (the case $w_{2}\in K^{\times}[0,1]^{T}$ is similar). Then $\mathrm{Stab}_{G}(w_{2})=\langle ac^{2}\rangle$, and $w_{3}$ can not be a scalar multiple of $0,1]^{T}$ (since the stabilizer of $0,1]^{T}\in W_{3}$ is also $\langle ac^{2}\rangle$). This means that $w_{3}$ is a non-zero scalar multiple of $1,0]^{T}$ as well, and so none of $\{p_{yyz},p_{yzz},p^{-}_{yyz},p^{-}_{yzz},p^{-}_{zzz}\}$ vanishes on $v=(w_{2},w_{3})$. These are relative invariants of degree $3$, and their weights represent all the five non-trivial characters.
Case 1.b: $q_{yy}(w_{2})=0=p^{-}_{zzz}(w_{3})$. Then $w_{2}$ is a non-zero scalar multiple of $1,0]^{T}$ or $0,1]^{T}$, and $w_{3}$ is a non-zero scalar multiple of $1,1]^{T}$, $1,\omega]^{T}$, $1,\omega^{2}]^{T}$. It follows that none of $\{p_{yyz},p_{yzz},p^{-}_{yyz},p^{-}_{yzz},p^{-}_{yyy}\}$ vanishes on $v$. These are relative invariants of degree $3$, and their weights represent all the five non-trivial characters.
@@
Case 1.d: $p^{-}_{yyy}(w_{2})=0=p^{-}_{zzz}(w_{3})$. Then both of $w_{2},w_{3}$ are non-zero scalar multiples of $1,1]^{T}$, $1,\omega]^{T}$, or $1,\omega^{2}]^{T}$. Moreover, $w_{2}$ and $w_{3}$ are linearly independent (otherwise they would have the same stabilizer). Thus none of $\{q_{yy},q_{zz},q^{-}_{yz},q_{yy}q^{-}_{yz},q_{zz}q^{-}_{yz}\}$ vanishes on $v$. These are relative invariants of degree at most $4$, and their weights represent all the five non-trivial characters.